# Evaluating Decision Rules Across Many Weak Experiments

### v1

@@
Abstract
Technology firms conduct randomized controlled experiments (“A/B tests”) to learn which actions to take to improve business outcomes. In firms with mature experimentation platforms, experimentation programs can consist of many thousands of tests. To scale experimentation effectively, firms rely on decision rules: standard operating procedures for mapping the results of an experiment to a choice of treatment arm to launch to the general user population. Despite the critical role of decision rules in translating experimentation into business decisions, rigorous guidance on how to evaluate and choose decision rules is scarce. This paper proposes to evaluate decision rules based on their cumulative returns to business north star metrics. Although this quantity is intuitive and easy to explain to decision-makers, estimating it can be non-trivial, especially when experiments have weak signal-to-noise ratios. We develop a cross-validation estimator that is much less biased than the naive plug-in estimator under conditions realistic to digital experimentation. We demonstrate the efficacy of our approach via a case study of 123 historical A/B tests at Netflix, where we used it to show that a new decision rule would increase cumulative returns to the north star metric by an estimated $33\%$, leading directly to the adoption of the new rule.
Keywords: A/B testing, experimentation, decision rules, proxy metrics, meta-analysis
1 Introduction
Online randomized controlled experimentation (aka A/B testing) is the workhorse of product innovation in many technology companies (Deng et al., 2017; Thomke, 2020). Netflix is no exception to this rule, running thousands of A/B tests per year on many diverse parts of the business (Ejdemyr et al., 2024). This scale and velocity necessitate the adoption of norms and practices that are easily communicated to and by scientists, engineers, and product managers alike (Bojinov et al., 2025).
In particular, companies commonly rely on decision rules: standard operating procedures that map the outcomes of an experiment to a “winning” treatment arm to launch to the general user population. Letting $Y$ denote a north star business reward, such as long-term user retention or revenue, a simple example of a decision rule is:
Launch the treatment arm with the largest effect on $Y$. (1)
@@ 1 Introduction
  • Guardrail metrics, which only permit launches if they do not show a decline in a metric that is negatively correlated with $Y$, or which is important to minimize for other reasons (e.g. app crashes, customer service contacts) (Deng and Shi, 2016).
@@
How should organizations evaluate and choose between these (and many other possible) decision rules? We propose evaluating decision rules based on their cumulative returns, which are not measurable in any one experiment but can be estimated across many historical experiments (Ejdemyr et al., 2024). Intuitively, our method evaluates a decision rule by asking, “What would the cumulative impact on $Y$ have been if all experiments were decided using this rule?”
@@
Although this quantity is intuitive and easy to communicate to decision-makers, simple backward-looking estimates of cumulative returns can be severely biased due to a winner’s curse: since winning arms are sometimes chosen due to noise, the underlying treatment effects of winning arms are likely smaller than the raw historical lifts suggest (Lee and Shen, 2018; Ejdemyr et al., 2024). A consequence of the winner’s curse is that rules that exploit spurious correlations can seemingly perform well in past experiments, but generalize poorly to future experiments. This bias persists even with an infinite number of experiments so long as each experiment has a finite sample size, and is exacerbated when experiments have a low signal-to-noise ratio, as is endemic to digital experimentation (Larsen et al., 2024; Kohavi et al., 2014).
@@
To overcome the winner’s curse, we demonstrate theoretically and empirically that a cross-validation estimator using experiment-splitting (cf. Coey and Cunningham, 2019) is many times less biased than the plug-in estimator under conditions realistic to digital experimentation. The cross-validation estimator separates the data used to choose the winning arm from the data used to evaluate its returns, eliminating the winner’s curse bias. The tradeoff is a small negative bias that is easily mitigated in practice.
@@
Our work is closely related to research on optimizing the returns to industrial-scale experimentation programs (Azevedo et al., 2018, 2023; Sudijono et al., 2024). This literature highlights how conventions borrowed from scientific practice (e.g. launching if and only if $p<0.05$) generally fail to maximize north star business metrics. For example, if the potential impact of innovations is heavy-tailed, firms should often run many ex ante underpowered experiments (Azevedo et al., 2018) and ship based on much more relaxed p-value thresholds (Sudijono et al., 2024). On the other hand, a stricter p-value threshold may be preferable for firms that associate a high cost with false discoveries (Kohavi and Chen, 2024). Our work calls for addressing these questions empirically via the evaluation of potential decision rules over a fixed set of past experiments. In our experience, evidence of this kind is essential for convincing decision-makers to consider different rules, as opposed more abstract arguments based on a superpopulation of hypothetical experiments.
@@
Our work also relates to the literature on proxy metrics in experimentation. Prior work has taken two distinct approaches to constructing proxy metrics. The first approach assesses the degree to which the proxy and north star metrics agree or disagree in sign, possibly accounting for statistical significance (Richardson et al., 2023; Tripuraneni et al., 2024; Jeunen and Ustimenko, 2024). This reflects two desirable properties of proxy metrics, namely directional alignment and statistical sensitivity, with less emphasis on an underlying structural model. The second approach targets specific structural estimands based on the north star metric: Bibaut et al. (2024) develops estimators for the covariance in true treatment effects between the proxy and north star (as opposed to the covariance of estimated treatment effects, which is subject to spurious finite-sample correlations), while Athey et al. (2019) proposes to construct an index of intermediate outcomes to target the unobserved lift in the north star metric. Taking a more general perspective, this paper treats proxy metrics as a special case of decision rules, although our methodology is also robust to spuriously correlated noise.
@@
  • Section 2 outlines a framework for evaluating decision rules based on their cumulative returns. Our framework is suitable for many practical applications including choosing p-value thresholds (Kohavi and Chen, 2024; Sudijono et al., 2024) and selecting proxy metrics or surrogate models (Athey et al., 2019; Bibaut et al., 2024; Richardson et al., 2023; Tripuraneni et al., 2024; Jeunen and Ustimenko, 2024; Kharitonov et al., 2017).
  • Section 3 provides an estimator for cumulative returns using past A/B tests, and specifies the conditions under which our estimator consistently selects the best rule from a finite set of candidates as the number of experiments (not the number of units in each experiment) increases.
  • Section 4 illustrates the benefits of our framework in a simplified example. We use our method to evaluate and select between candidate proxy metrics in simulated data and demonstrate the favorable properties of our approach.
  • Lastly, Section 5 describes the real-world application of our framework to select decision rules at Netflix. Using our method, we demonstrated that a new decision rule would increase the cumulative returns on a sample of past experiments by approximately 33%. This evidence was instrumental to the adoption of our rule for deciding new experiments at Netflix.
2 Framework and Notation
We study a setting in which a firm has conducted $N$ past experiments indexed by $i=1,\ldots,N$. Experiment $i$ has $K_{i}$ arms, indexed by $k=1,\ldots,K_{i}$, with each arm consisting of $M_{i}$ units, indexed by $m=1,\ldots,M$. We typically think of $M_{i}$ as large in online experiments, numbering in the thousands or even millions. However, $M_{i}$ is also bounded even for the largest firms, which have a finite user base to allocate across many parallel experiments (Tang et al., 2010).
@@ 2 Framework and Notation
We assume there is a known reward function $\psi:\mathbb{R}^{J}\to\mathbb{R}$, and let
@@
be the mean reward from arm $k$ in experiment $i$. For example, if of the $J$ metrics we treat the first one as the reward, we set $\psi(o)=o_{1}$. The other $J-1$ metrics may still be used (or not) for making a decision. [^1] For example, the second metric may be a proxy with lower variance whose mean might have some bias to the true reward. Or, the other metrics are pre-treatment variables we can use for variance reduction via post-stratification or regression adjustment.
@@
Lastly, it is also common to only consider launching arms for which $\phi$ is statistically significantly greater than a reference arm. Denoting the reference arm as $k=1$ and the set of such arms as $\mathcal{S}(\mathbf{O}^{i})$, this rule is written as:
@@
We can envision many other variants and extensions, including correcting for pre-intervention variables or using lower confidence bounds. In any event, given the many reasonable-sounding alternatives, we want to avoid choosing rules arbitrarily, and instead choose between them in a data-driven manner.
3 Proposed Methodology
@@
$\hat{R}_{i}^{naive}(D)=\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\sum_{k=1}^{K_{i}}\mathbf{1}(D(\mathbf{O}^{i})=k)\psi(O^{i}_{km}).$ (2)
@@
$\mathbb{E}[\hat{R}_{i}^{naive}]-R_{iD(\mathbf{O}^{i})}$ (3)
@@
Our proposed cross-validation estimator uses experiment splitting to eliminate the winner’s curse:
{mdframed}
[linewidth=0.3mm, linecolor=black, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt] Cross-Validation Algorithm:
  • Randomly split the units in each experiment $i$ into $P_{i}$ folds.
  • For $i=1,\ldots,N$ and $p=1,\ldots,P_{i}$:
    • Apply the decision rule $D(\mathbf{O}_{-p}^{i})$ to select a winning arm.
    • Estimate the reward of $D(\mathbf{O}_{-p}^{i})$ on the $p$th fold as:
      $\hat{R}^{CV}_{ip}=\frac{1}{M_{ip}}\sum_{m=1}^{M_{ip}}\sum_{k=1}^{K_{i}}\mathbf{1}(D(\mathbf{O}_{-p}^{i})=k)\psi(O^{i}_{km})$
      where $M_{ip}$ is the number of units in the $p$th fold for experiment $i$.
  • Lastly, estimate the (weighted) mean reward as the mean of $\hat{R}^{CV}_{ip}$ over all folds and experiments:
    $\hat{\Gamma}^{CV}$ $=$ $\frac{1}{\sum_{i=1}^{N}P_{i}}\sum_{i=1}^{N}\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}$
In a nutshell, our proposed estimator separates the data used to choose the winning arm from the data used to estimate the reward of that arm, eliminating the winner’s curse. However, the cross-validation estimator is also biased because the decision rule is no longer applied to the full sample, changing the probability with which each arm is chosen:
$\mathbb{E}[\hat{R}^{CV}_{ip}]-R_{iD(\mathbf{O}^{i})}=$ (4)
@@
First, we can increase the number of folds $P$. While the CV estimator is biased for the expected reward of the rule applied to the full data, when all folds are of the same size $M_{ip}=M_{i}/P$ then it is unbiased for the expected reward of the rule applied to the first $(P-1)/P$ fraction of the data. These two estimands will become very close as the number of folds $P$ approaches the sample size $M_{i}$. However, at best we can use leave-one-out cross-validation to test the rule that uses $M_{i}-1$ observations, which is still technically different than using all $M_{i}$ observations. While this difference should be very small in applications, it can be meaningful when experiments are just barely powered to detect effects between arms. As a practical matter, we recommend assessing the sensitivity of the CV estimator to the choice of $P$.
Second, we can provide theoretical guarantees in the specific setting where the amount of data in each experiment is itself random. We show that, up to a scaling, cross-validation incurs zero bias on average in this realistic setting. Specifically, suppose that $M_{i}$ is random and Poisson-distributed. This is consistent, for example, with enrolling a fixed percentage of all incoming units into each experiment over a fixed time period. Consider a $-\ell$-out setting where there are $P_{i}=\prod_{j=0}^{\ell-1}(M_{i}-j)$ folds consisting of all the $M_{ip}=\ell$ sized subsets of the units $\{1,\dots,M_{i}\}$.
@@
Suppose $M_{i}\sim\mathrm{Poisson}(M_{0})$, then for the $-\ell$-out CV estimator, we have $\mathbb{E}[M_{0}^{-\ell}\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}]=\mathbb{E}[R_{iD(\mathbf{O}^{i})}]$.
@@
Define $R_{i}(t)=\mathbb{E}[R_{iD(\mathbf{O}^{i})}\mid M_{i}=t]$. Then $\mathbb{E}[\hat{R}^{CV}_{ip}\mid M_{i}]=R_{i}(M_{i}-\ell)$, so $\mathbb{E}[\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}\mid M_{i}]=P_{i}R_{i}(M_{i}-\ell)$. On the other hand, $\mathbb{E}[R_{iD(\mathbf{O}^{i})}]=\mathbb{E}[R_{i}(M_{i})]$ by total expectation. The Stein-Chen identity (Chen, 1975) states that for $X\sim\mathrm{Poisson}(\lambda)$, we have $\lambda\mathbb{E}[f(X)]=\mathbb{E}[Xf(X-1)]$ for any $f:\mathbb{Z}\to\mathbb{R}$. Applying this identity $\ell$ times we obtain $\lambda^{\ell}\mathbb{E}[f(X)]=\mathbb{E}[X(X-1)\cdots(X-\ell+1)f(X-\ell)]$. Since $P_{i}=\prod_{j=0}^{\ell-1}(M_{i}-j)$ we obtain the conclusion. ∎
The practical implication of Theorem 3.1 is that, if $M_{i}$ is Poisson, the cross-validation estimator will be able to select the best rule (in terms of $R$) from a finite set of decision rules as the number of experiments $N$ goes to infinity. Importantly, this does not require the size of each experimental arm $M_{i}$ to diverge. In general, letting $M_{i}\to\infty$ is an unrealistic asymptotic regime because all firms have a finite user pool from which to allocate users.
@@
Suppose $M_{i}\sim\mathrm{Poisson}(M_{0})$ and $\|\psi(O_{km}^{i})\|_{\infty}<\infty$. Fix $\left\lvert\mathcal{D}\right\rvert<\infty$ and let
$\hat{D}\in\operatornamewithlimits{argmax}_{D\in\mathcal{D}}\hat{\Gamma}^{CV}(D).$
Then $\max_{D\in\mathcal{D}}\Gamma(D)-\Gamma(\hat{D})=O_{p}(\log\left\lvert\mathcal{D}\right\rvert/\sqrt{NM_{0}})$.
@@
For simplicity, let us focus on the case of leave-one-out ($P_{i}=M_{i}$). Let $X_{i}(D)=M_{0}^{-1}\sum_{p=1}^{P_{i}}\hat{R}_{ip}^{CV}(D)$ and observe that $\hat{D}\in\operatornamewithlimits{argmax}_{D\in\mathcal{D}}\frac{1}{N}\sum_{i=1}^{N}X_{i}(D)$. Therefore $\max_{D\in\mathcal{D}}\Gamma(D)-\Gamma(\hat{D})\leq 2\sup_{D\in\mathcal{D}}\left|\frac{1}{N}\sum_{i=1}^{N}\bar{X}_{i}(D)\right|$ where $\bar{X}_{i}(D)=X_{i}(D)-R_{iD(\mathbf{O}^{i})}$. Theorem 3.1 established that $\mathbb{E}\bar{X}_{i}(D)=0$. Let $C=\|\psi(O_{km}^{i})\|_{\infty}<\infty$. Then
$\textstyle\mathbb{E}\exp(\frac{t}{N}\sum_{i=1}^{N}\bar{X}_{i}(D))$ $\textstyle\leq\prod_{i=1}^{N}\mathbb{E}\exp(t(M_{i}/M_{0}+1)C/N)$
$=\exp(NM_{0}(e^{tC/(NM_{0})}-1+tC/(NM_{0})))$
$\leq\exp(t^{2}C^{2}/(NM_{0})),$
for $tC\leq NM_{0}$. The rest follows by a Chernoff and union bound. ∎
4 Example: Choosing a Proxy Metric from Past Experiments
Above, we developed notation to formalize a target estimand (expected reward), demonstrated that an estimator based on cross-validation does not suffer from the winner’s curse, and discussed ways to counteract the remaining bias in the CV estimator. This section builds further intuition by focusing on a specific application of our method: choosing a good proxy metric for deciding experiments (Bibaut et al., 2024; Tripuraneni et al., 2024; Richardson et al., 2023).
Proxy metrics are commonly used in digital experimentation to accelerate decision-making by replacing statistically insensitive and/or long-term metrics (e.g., user retention) with more sensitive and/or short-term metrics (e.g. user engagement). However, choosing a good proxy metric is nontrivial as interventions that increase the proxy metric may not increase the north star metric, even if the two variables are strongly correlated in either experimental (Bibaut et al., 2024) or observational (Athey et al., 2019) data. To decide between different proxy metrics, then, we can apply the above estimators for the expected reward associated with each metric and choose the metric(s) which lead to the largest estimated reward.
4.1 Setup
To simplify our illustration, we assume $N$ past experiments with $K=2$ arms each and $M_{i}=M$ units per arm. In each experiment, we estimate treatment effects $\hat{\tau}_{Y}^{i}$ and $\hat{\tau}_{S}^{i}$ on the north star and proxy metric, respectively. We assume that $M$ is large and that our treatment effect estimator is the difference in sample means, so that the observed treatment effects are Gaussian conditional on the true treatment effects:
$(\hat{\tau}_{Y}^{i},\hat{\tau}_{S}^{i})\sim\mathcal{N}\left((\tau_{Y}^{i},\tau_{S}^{i}),\frac{2}{M}\Omega\right),$ (5)

### v2

@@
Abstract.
Technology firms conduct randomized controlled experiments (“A/B tests”) to learn which actions to take to improve business outcomes. In firms with mature experimentation platforms, experimentation programs can consist of many thousands of tests. To effectively scale experimentation, firms rely on decision rules: standard operating procedures for mapping the results of an experiment to a choice of treatment arm to launch to the general user population. Despite the critical role of decision rules in translating experimentation into business decisions, rigorous guidance on how to evaluate and choose decision rules is scarce. This paper proposes to evaluate decision rules based on their cumulative returns to business north star metrics. Although intuitive and easy to explain to decision-makers, this quantity can be difficult to estimate, especially when experiments have weak signal-to-noise ratios. We develop a cross-validation estimator that is much less biased than the naive plug-in estimator under conditions realistic to digital experimentation. We demonstrate the efficacy of our approach via a case study of 123 historical A/B tests at Netflix, where we used it to show that a new decision rule would have increased cumulative returns to the north star metric by an estimated $33\%$, directly leading to the adoption of the new rule.
A/B testing, experimentation, decision rules, proxy metrics, meta-analysis
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737217isbn: 979-8-4007-1454-2/2025/08ccs: Mathematics of computing Probability and statisticsccs: Computing methodologies Rule learningccs: Computing methodologies Cross-validation
1. Introduction
Online randomized controlled experimentation (aka A/B testing) is the workhorse of product innovation in many technology companies (Deng et al., 2017; Thomke, 2020). Netflix is no exception to this rule, running thousands of A/B tests per year on many diverse parts of the business (Ejdemyr et al., 2024). This scale necessitates the adoption of norms and practices that are easily communicated between scientists, engineers, and product managers alike (Bojinov et al., 2025).
In particular, companies often rely on decision rules: standard operating procedures that map the outcomes of an experiment to a “winning" treatment arm to launch to the general user population. Letting $Y$ denote a north star business reward, such as long-term user retention or revenue, a simple example of a decision rule is:
(1) Launch the treatment arm with the largest effect on $Y$.
@@ 1. Introduction
  • Guardrail metrics, which only permit launches if they do not show a decline in a metric that is negatively correlated with $Y$, or which is important to minimize for other reasons (e.g., app crashes, customer service contacts) (Deng and Shi, 2016).
@@
How should organizations evaluate and choose between these (and many other possible) decision rules? We propose evaluating decision rules based on their cumulative returns, which are not measurable in any one experiment, but can be estimated across many historical experiments (Ejdemyr et al., 2024). Intuitively, our method evaluates a decision rule by asking, “What would the cumulative impact on $Y$ have been if all experiments were decided using this rule?”
@@
Although this quantity is intuitive and easy to explain to decision-makers, simple backward-looking estimates of cumulative returns can be severely biased due to a winner’s curse: since winning arms are sometimes chosen due to noise, the underlying treatment effects of winning arms are likely smaller than the raw historical lifts suggest (Lee and Shen, 2018; Ejdemyr et al., 2024). A consequence of the winner’s curse is that rules that exploit spurious correlations can seemingly perform well in past experiments, but generalize poorly to future experiments. This bias persists with even an infinite number of experiments so long as each experiment has a finite sample size, and is exacerbated by the low signal-to-noise ratios endemic to digital experimentation (Larsen et al., 2024; Kohavi et al., 2014).
@@
To overcome the winner’s curse, we demonstrate theoretically and empirically that a cross-validation estimator using experiment splitting (Coey and Cunningham, 2019) is many times less biased than the plug-in estimator under conditions realistic to digital experimentation. The cross-validation estimator separates the data used to choose the winning arm from the data used to evaluate its returns, eliminating the winner’s curse. The tradeoff is a small negative bias that is easily mitigated in practice.
@@
Our work is closely related to research on optimizing the returns to industrial-scale experimentation (Azevedo et al., 2018, 2023; Sudijono et al., 2024). This literature highlights how conventions borrowed from scientific practice (e.g., launching if and only if $p<0.05$) generally fail to maximize north star business metrics. For example, if the potential impact of innovations is heavy-tailed, firms should run many ex ante underpowered experiments (Azevedo et al., 2018) and launch based on much more relaxed p-value thresholds (Sudijono et al., 2024). On the other hand, a stricter $p$-value threshold may be preferable for firms that associate a high cost with false discoveries (Kohavi and Chen, 2024). Our work calls for addressing these questions empirically via the evaluation of decision rule candidates across a fixed set of past experiments. In our experience, evidence of this kind is essential to convincing decision-makers to consider different rules, as opposed to more abstract arguments based on a superpopulation of hypothetical experiments.
@@
Our work also relates to the literature on proxy metrics in experimentation. Previous work has taken two distinct approaches to constructing proxy metrics. The first approach assesses the degree to which the proxy and north star metrics agree or disagree in sign, possibly accounting for statistical significance (Richardson et al., 2023; Tripuraneni et al., 2024; Jeunen and Ustimenko, 2024). This reflects two desirable properties of proxy metrics, namely directional alignment and statistical sensitivity, with less emphasis on an underlying structural model. The second approach targets specific structural estimands based on the north star metric: (Bibaut et al., 2024) develops estimators for the covariance in true treatment effects between the proxy and north star (as opposed to the covariance of estimated treatment effects, which is subject to spurious finite-sample correlations), while (Athey et al., 2019) proposes to construct an index of intermediate outcomes to target the unobserved lift in the north star metric. Taking a more general perspective, we treat proxy metrics as a special case of decision rule and propose to estimate their performance empirically, although our methodology also ensures robustness to spuriously correlated noise.
@@
  • Section 2 outlines a framework for evaluating decision rules based on their cumulative returns. Our framework is suited to many practical applications, including the choice of $p$-value thresholds (Kohavi and Chen, 2024; Sudijono et al., 2024) and proxy metrics or surrogate models (Athey et al., 2019; Bibaut et al., 2024; Richardson et al., 2023; Tripuraneni et al., 2024; Jeunen and Ustimenko, 2024; Kharitonov et al., 2017).
  • Section 3 provides an estimator for cumulative returns using past A/B tests and specifies the conditions under which our estimator consistently selects the best rule from a finite set of candidates as the number of experiments (not the number of units in each experiment) increases.
  • Section 4 illustrates the benefits of our framework in a stylized example. We use our method to evaluate and select between candidate proxy metrics in simulated data and demonstrate the favorable properties of our approach.
  • Lastly, Section 5 describes the real-world application of our framework to select decision rules at Netflix. Using our method, we demonstrated that a new decision rule would increase the cumulative returns on a sample of past experiments by approximately 33%. This evidence led directly to the adoption of our rule to decide new experiments at Netflix.
2. Framework and Notation
We study a setting in which a firm has conducted $N$ past experiments indexed by $i=1,\ldots,N$. Experiment $i$ has $K_{i}$ arms, indexed by $k=1,\ldots,K_{i}$, with each arm consisting of $M_{i}$ units, indexed by $m=1,\ldots,M$. We typically think of $M_{i}$ as large in online experiments, numbering in the thousands or even millions. However, $M_{i}$ is also bounded for even the largest firms, which have a finite user base to allocate across many parallel experiments (Tang et al., 2010). Therefore, as in (Bibaut et al., 2024), our theoretical results focus on the asymptotic regime in which the number of experiments, not the number of users, goes to infinity.
@@ 2. Framework and Notation
We assume there is a known reward function $\psi:\mathbb{R}^{J}\to\mathbb{R}$, and let:
@@
be the mean reward from arm $k$ in experiment $i$. For example, if of the $J$ metrics we treat the first one as the reward, we set $\psi(o)=o_{1}$. The other $J-1$ metrics may still be used (or not) for making a decision. [^1] For example, the second metric may be a proxy with lower variance whose mean might have some bias to the true reward. Or, the other metrics may be pretreatment variables used for variance reduction via post-stratification or regression adjustment.
@@
Lastly, it is also common to consider only launching arms for which $\phi$ is statistically significantly greater than a reference arm. Denoting the reference arm as $k=1$ and the set of such arms as $\mathcal{S}(\mathbf{O}^{i})$, this rule is written as:
@@
We can envision many other variants and extensions, including correcting for pretreatment variables or using lower confidence bounds. In any event, given the many reasonable-sounding alternatives, we want to avoid choosing rules arbitrarily, and instead choose between them in a data-driven manner.
3. Proposed Methodology
@@
(2) $\hat{R}_{i}^{naive}(D)=\frac{1}{M_{i}}\sum_{m=1}^{M_{i}}\sum_{k=1}^{K_{i}}\mathbf{1}(D(\mathbf{O}^{i})=k)\psi(O^{i}_{km}).$
@@
(3) $\mathbb{E}[\hat{R}_{i}^{naive}]-R_{iD(\mathbf{O}^{i})}$
@@
Alternatively, our proposed cross-validation estimator uses experiment splitting to eliminate the winner’s curse:
Cross-Validation Algorithm Randomly split the units in experiment $i$ into $P_{i}$ folds. For $i=1,\ldots,N$ and $p=1,\ldots,P_{i}$: Apply the decision rule $D(\mathbf{O}_{-p}^{i})$ to select a winning arm. Estimate the reward of $D(\mathbf{O}_{-p}^{i})$ on the $p$th fold as: $\hat{R}^{CV}_{ip}=\frac{1}{M_{ip}}\sum_{m=1}^{M_{ip}}\sum_{k=1}^{K_{i}}\mathbf{1}(D(\mathbf{O}_{-p}^{i})=k)\psi(O^{i}_{km})$ where $M_{ip}$ is the number of units in fold $p$ for experiment $i$. Lastly, estimate the mean reward as the (weighted) mean of $\hat{R}^{CV}_{ip}$ over all folds and experiments: $\hat{\Gamma}^{CV}$ $=$ $\frac{1}{\sum_{i=1}^{N}P_{i}}\sum_{i=1}^{N}\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}$
In a nutshell, our estimator separates the data used to choose the winning arm from the data used to estimate the reward of that arm, eliminating the winner’s curse. Note, however, that the cross-validation estimator is also biased because the decision rule is no longer applied to the full sample, changing the probability with which each arm is chosen:
(4) $\mathbb{E}[\hat{R}^{CV}_{ip}]-R_{iD(\mathbf{O}^{i})}=$
@@
First, we can increase the number of folds $P$. While the CV estimator is biased for the expected reward of the rule applied to the full data, when all folds are of the same size $M_{ip}=M_{i}/P$ then it is unbiased for the expected reward of the rule applied to the first $(P-1)/P$ fraction of the data. These two estimands will become very close as the number of folds $P$ approaches the sample size $M_{i}$. However, in the best case we can use leave-one-out cross-validation to test the rule that uses $M_{i}-1$ observations, which is still technically different from using all $M_{i}$ observations. Although this difference should be very small in applications, it can be meaningful when experiments are just barely powered to detect effects between arms. In practice, we recommend assessing the sensitivity of the CV estimator to the choice of $P$.
Second, we can provide theoretical guarantees in the specific setting where the amount of data in each experiment is itself random. We show that, up to a scaling, cross-validation incurs zero bias on average in this realistic setting. Specifically, suppose that $M_{i}$ is random and Poisson-distributed. This is consistent, for example, with enrolling a fixed percentage of all incoming units into each experiment over a fixed time period. Consider $-\ell$-out cross-validation, where there are $P_{i}=\ell!^{-1}\prod_{j=0}^{\ell-1}(M_{i}-j)$ folds consisting of all subsets of $M_{ip}=\ell$ units $\{1,\dots,M_{i}\}$.
@@
Suppose $M_{i}\sim\mathrm{Poisson}(M_{0})$. Then, for the $-\ell$-out CV estimator, we have $\mathbb{E}[M_{0}^{-\ell}\ell!\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}]=\mathbb{E}[R_{iD(\mathbf{O}^{i})}]$.
@@
Define $R_{i}(t)=\mathbb{E}[R_{iD(\mathbf{O}^{i})}\mid M_{i}=t]$. Then $\mathbb{E}[\hat{R}^{CV}_{ip}\mid M_{i}]=R_{i}(M_{i}-\ell)$, so $\mathbb{E}[\sum_{p=1}^{P_{i}}\hat{R}^{CV}_{ip}\mid M_{i}]=P_{i}R_{i}(M_{i}-\ell)$. On the other hand, $\mathbb{E}[R_{iD(\mathbf{O}^{i})}]=\mathbb{E}[R_{i}(M_{i})]$ by total expectation. The Stein-Chen identity (Chen, 1975) states that, for $X\sim\mathrm{Poisson}(\lambda)$, we have $\lambda\mathbb{E}[f(X)]=\mathbb{E}[Xf(X-1)]$ for any $f:\mathbb{Z}\to\mathbb{R}$. Applying this identity $\ell$ times we obtain $\lambda^{\ell}\mathbb{E}[f(X)]=\mathbb{E}[X(X-1)\cdots(X-\ell+1)f(X-\ell)]$. Since $P_{i}=\ell!^{-1}\prod_{j=0}^{\ell-1}(M_{i}-j)$, we obtain the conclusion. ∎
The practical implication of Theorem 3.1 is that, if $M_{i}$ is Poisson, the cross-validation estimator will be able to select the best rule (in terms of $R$) from a finite set of decision rules as the number of experiments $N$ goes to infinity. Importantly, this does not require the size of each experimental arm $M_{i}\to\infty$, which is an unrealistic asymptotic regime given that all firms have a finite user pool to allocate to many parallel experiments (Tang et al., 2010). We formalize and prove the result for leave-one-out cross-validation ($P_{i}=M_{i}$) below.
@@
Suppose $M_{i}\sim\mathrm{Poisson}(M_{0})$ and $\|\psi(O_{km}^{i})\|_{\infty}<\infty$. Fix $\left\lvert\mathcal{D}\right\rvert<\infty$ and let:
$\hat{D}\in\operatornamewithlimits{argmax}_{D\in\mathcal{D}}\hat{\Gamma}^{CV}(D)$
with $\ell=1$. Then $\max_{D\in\mathcal{D}}\Gamma(D)-\Gamma(\hat{D})=O_{p}\left(\sqrt{\log|\mathcal{D}|/(NM_{0})}\right)$.
@@
Let $X_{i}(D)=M_{0}^{-1}\sum_{m=1}^{M_{i}}\hat{R}_{ip}^{CV}(D)$ and observe that $\hat{D}\in\operatornamewithlimits{argmax}_{D\in\mathcal{D}}\frac{1}{N}\sum_{i=1}^{N}X_{i}(D)$. Therefore, $\max_{D\in\mathcal{D}}\Gamma(D)-\Gamma(\hat{D})\leq 2\sup_{D\in\mathcal{D}}\left|\frac{1}{N}\sum_{i=1}^{N}\bar{X}_{i}(D)\right|$ where $\bar{X}_{i}(D)=X_{i}(D)-R_{iD(\mathbf{O}^{i})}$. Theorem 3.1 established that $\mathbb{E}\bar{X}_{i}(D)=0$. Let $C=\|\psi(O_{km}^{i})\|_{\infty}<\infty$. Then:
$\mathbb{E}\left[\exp\left(\frac{t}{N}\sum_{i=1}^{N}\bar{X}_{i}(D)\right)\right]$
$\quad=\left(\mathbb{E}\left[\exp\left(\frac{t}{N}\bar{X}_{i}(D)\right)\right]\right)^{N}$
$\quad=\left(e^{-s\Gamma(D)}\mathbb{E}\left[\exp\left(sX_{i}(D)\right)\right]\right)^{N}\bigg{|}_{s=t/N}$
$\quad=\left(e^{-s\Gamma(D)}\exp\left(M_{0}\left(\mathbb{E}\left[e^{(s/M_{0})\hat{R}_{ip}^{CV}}\right]-1\right)\right)\right)^{N}\bigg{|}_{s=t/N}$
$\quad\leq\left(\exp\left(\frac{s^{2}\mathbb{E}[(\hat{R}_{ip}^{CV})^{2}]}{M_{0}}\right)\right)^{N}\bigg{|}_{s=t/N}$
$\quad\leq\exp\left(\frac{t^{2}C^{2}}{NM_{0}}\right)$
for $|s|C\leq M_{0}$. The rest follows by a Chernoff and union bound. ∎
4. Example: Choosing a Proxy Metric from Past Experiments
Above, we developed notation to formalize a target estimand (expected reward), demonstrated that an estimator based on cross-validation does not suffer from the winner’s curse, and discussed ways to counteract the remaining bias in the CV estimator. This section further builds intuition by focusing on a specific application of our method: choosing a good proxy metric for deciding experiments (Bibaut et al., 2024; Tripuraneni et al., 2024; Richardson et al., 2023).
Proxy metrics are commonly used in digital experimentation to accelerate decision-making by replacing statistically insensitive and/or long-term metrics (e.g., user retention) with more sensitive and/or short-term metrics (e.g., user engagement). However, choosing a good proxy metric is non-trivial as interventions that increase the proxy metric may not increase the north star metric, even if the two variables are strongly correlated in either experimental (Bibaut et al., 2024) or observational (Athey et al., 2019) data. To decide between different proxy metrics, we can apply the above estimators for the expected reward associated with each proxy and choose the one(s) which leads to the largest estimated reward.
4.1. Setup
To simplify our numerical example, we assume $N$ past experiments with $K=2$ arms each and $M_{i}=M$ units per arm. In each experiment, we estimate treatment effects $\hat{\tau}_{Y}^{i}$ and $\hat{\tau}_{S}^{i}$ on the north star and proxy metric, respectively. We assume that $M$ is large and that our treatment effect estimator is the difference in sample means, so that the observed treatment effects are Gaussian conditional on the true treatment effects:
(5) $(\hat{\tau}_{Y}^{i},\hat{\tau}_{S}^{i})\sim\mathcal{N}\left((\tau_{Y}^{i},\tau_{S}^{i}),\frac{2}{M}\Omega\right),$